One component of an ideal binary solution has a pure component vapor pressure equal to one-fourth the total pressure of the vapor mixture. The component constitutes 20 mole % of liquid phase. What is its mole fraction in the vapor phase?

1
Expert's answer
2022-02-08T21:04:02-0500

Using Raoult's law and Dalton law


"x_i=\frac{y_i\>p_{TOTAL}}{p_i^*}"


Where "x_i=" Mole fraction of "i" component in solution

"y_i=" Mole fraction of component "i" in gas phase

"p_{TOTAL}=" Total vapor pressure

"p_i^*=" Equilibrium vapor pressure of the pure component "i"


"\frac{20}{100}=\frac{y_i\>p_{TOTAL}}{\frac{1}{4}\>p_{TOTAL}}"


"y_i=\frac{1}{4}×\frac{1}{5}"


"=\frac{1}{20}"